2 1/2 x - 2 1/4 = 3 2/5 ?
Answer by checkley77(12844)

2 1/2 x - 2 1/4 = 3 2/5
2.5x-2.25=3.4
2.5x=3.4+2.25
2.5x=5.65
x=5.65/2.5
x=2.26 ans.